You are here: Foswiki>FLAO Web>TowerLab>LabTest20100505 (06 May 2010, AlessandroGherardi)Edit Attach
-- AlessandroGherardi - 06 May 2010

Test slitte:

movimentazione ok

homing ok

Ripristino filterwheels:

Problema: Sensori di homing non alimentati

Soluzione: saldato filo sul connettore J2.72 box sx (con J23.21 board)

Test filterwheels:

movimentazione ok

homing ok
Topic revision: r1 - 06 May 2010, AlessandroGherardi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback